Your company is considering investing in Project X. Your analysts estimate that this project will increase your company’s net cash flows by $96246 in year 1, $141845 in year 2, and $203878 in year 3. For convenience, they treat these cash flows as occurring at the end of the respective years. Your analysts decided that the appropriate discount rate for this project is 7% per year, compounded annually. What is the present value of the cash flows expected from Project X?
Answer #1

To get present value of each cash flow , we need to discount them using formula: present value = future value /((1+discount rate)^no. Of years)

Year1 cash flow present value=$ 96246/((1+0.07)^1) = $89949.53

Year2 cash flow present value= $141845 /((1+0.07)^2) = $ 123892.92

Year3 cash flow present value = $203878 /((1+0.07)^3) =$ 166425.18

Present value of project x = total of present values of all 3 years cash flows = 89949.53 + 123892.92 + 166425.18 = $380267.63